HUBBLE SPACE TELESCOPE DAILY REPORT #2257 PERIOD COVERED: 0000Z (UTC) 11/23/98 - 0000Z (UTC) 11/24/98 Daily Status Report as of 328/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Ten Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The loss of lock documented in 2.1 and HSTAR 6782 occurred during the ninth iteration of this proposal. Otherwise, the observations were completed as planned, and no further anomalies were noted. 1.4 Completed WF/PC-2 6503 (The Enigma Lens Q2345+007: Early Assembly of Dark Matter?) The WF/PC-2 was used to make deep observations of the lensing double quasar system Q2345+007. The current evidence suggests that this double QSO results from the gravitational lensing effects of a large compact mass at z = 1.49. If so, this early formation of compact structure is not compatible with many forms of dark matter. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Four Sets of NIC/1/2/3 7961 (Flats: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the flat fields during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. These flats will provide a monitor for particulate comtamination and detector lateral position {from the coronagraphic spot and FDA vignetting}. The proposal completed with no reported problems. 1.6 Completed STIS/CCD 7634 (CCD Flat-Field Stability and Cosmetic Defect Fraction) The Space Telescope Imaging Spaectrograph (CCD) was used to investigate and monitor the STIS CCD imaging flat-field stability and cosmetic defects over monthly periods though out Cycle 7. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Four Sets of STIS/MA1 7950 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a dark calibration image. This is to establish if the detectors are performing nominally, that no internal movement of the MCP has taken place, and establish dark level in orbit for a given operating temperature. These darks will be used in later image reduction and also trended over instrument life time as indicators of detector degradation or damage. The observations were completed as planned, and no problems were reported. 1.8 Completed Five Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed Five Sets of STIS/CCD 8057 (STIS CCD Performance Monitor {continued}) The Space Telescope Imaging Spectrograph (CCD) was used to monitor to overall CCD performance. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed Eight Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.11 Completed WF/PC-2 6651 (WLR Distances from Accurate Multi-Color WF/PC-2 Photometry of A & B Supergiants in M-31 and M-33) The WF/PC-2 was used to make multi-color photometric observations of NGC-598 (a.k.a., M-33). This is to study A and B supergiant stars in M-31 and M-33 with a precision far exceeding that of existing ground-based estimates. The loss of lock documented in 2.1 and HSTAR 6782 occurred during this proposal. Otherwise, the observations were completed as planned, and no further problems were reported. 1.12 Completed STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. The observations were completed as scheduled, and no problems were reported. 1.13 Completed Two Sets of STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. The loss of lock documented in 2.1 and HSTAR 6782 occurred during the first iteration of this proposal. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.14 Completed NIC/3 7962 (NICMOS Warm-up Focus Monitoring) The Near Infrared Camera (Cameras Three) was used to make observations of the globular cluster and nebula of NGC-3603 as part of the NICMOS warm-up focus monitoring. This proposal is used to determine and monitor the optimal focus settings for NIC/2 and NIC/3 during the dewar warm-up period at the end of 1998. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed WF/PC-2 6669 (Imaging of Two Dwarf Galaxies in Tidal Tails) The WF/PC-2 was used to make observations of tidal trails of the two merging dwarf galaxies in NGC-4038. These two dwarf galaxies are part of the M-81 group, and recent models suggest that it is interacting with, but has not formed from, the H-1 surrounding M81. The observations were completed as planned, and no anomalies were reported. 1.16 Completed NIC/2 7226 (Search for Massive Jupiters) The Near Infrared Camera (Camera Two) was used to make a corographic search for massive planets around nearby, young main sequence stars. Because of the extreme youth of these objects, any low-mass brown dwarf and planetary companions will still be in a higher luminosity phase and thus easily detectable. A dark calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.17 Completed Three Sets of NIC/2/3 8063 (NICMOS Pure Parallel Sponge - Backup) The Near Infrared Camera (Cameras Two and Three) was used to make observations taken in parallel with the prime science observations. This proposal is PRE-the new FOM default. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 HSTAR 6782 describes a one-minute, 40-second loss of lock that occurred at 327/151057Z. The proposals mentioned in 1.3, 1.11, and 1.13 may have been affected when the take data flag went down. The acquisition at 328/013430Z defaulted to fine lock backup on FGS-1 only when the scan step limit was exceeded on FGS-3. HSTAR 6784 was written. Five proposals, all occurring during the period of the next report may have been affected: WF/PC-2 7344, NIC 7961, STIS 8062, NIC 7963, and STIS 8057. Scheduled Reacquisitions : 5 Successful : 5 2.2 FHST Updates: Scheduled : 25 Successful : 25 2.3 Operations Notes: The SSR EDAC error counter was cleared six times as directed by ROP SR-1. The NSSC-1 status buffer was dumped at 327/1642Z per ROP NS-3. HST telemetry was commanded to "P" format at 327/2013Z. Three TTRs were written: one for IP missing minor frames at 327/2147Z and at 327/2311Z, another for data inversions at 327/1444Z and 327/1521Z due to a STGT problem, and another for data hits at 328/0117Z and at 328/0839Z. A NICMOS EDAC single bit error occurred at 328/020755Z. The NICMOS internal error log was not dumped, but the ground system limit was raised to 48. The STIS MCE-2 reset at 328/091702Z while in a SAA interval and while the high voltage was off. HSTAR 6785 was written. The STIS flight software error counter was reset at 328/102436Z per ROP NS-12. Recovery will be performed via SMS commanding at 329/212854Z, the next time when the MAMA-2 high voltage will be turned on. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH